OpenStudy (anonymous):

A football is kicked off the flat ground at 23.0m/s at an angle of 30 degrees relative to the ground. Determine the total time it is in the air.

OpenStudy (anonymous):

A football is kicked off the flat ground at 23.0m/s at an angle of 30degrees relative to the ground. Determine the total time it is in the air. sin(30)=v_vert/23 30sin(23)=v_vert v_vert=11.5 v=u+at 0=11.5=(-9.8)xt 9.8(t)=11.5 t=11.5/9.8=1.173 Just copied and pasted that from Yahoo Answers, The answer goes more indepth...Refer here and make sure to check if its right: http://answers.yahoo.com/question/index?qid=20110929214236AAK6PZ7